OpenStudy (anonymous):

If m varies directly as p, and m=48 when p=8, find m when p is 4 If x varies inversely as y^2, and x=6 when y=10, find x when y=2 Help solving? I have a quiz soon so I need help on how to solve, please show work I do not want just the answers. It helps a lot if you draw it out for me so I can see it better. Thanks!!

OpenStudy (anonymous):

p = km, 8 = 48k, k = 1/6, 4 = 1/6m, m = 2/3

OpenStudy (anonymous):

how did you get from 4=1/6m to me= 2/3?

OpenStudy (anonymous):

m= **

OpenStudy (anonymous):

Hmm typo on my part, m = 24, I multiplied by the opposite of 1/6 which is 6/1 or 6

OpenStudy (anonymous):

yeah that's what I thought it should be haha

OpenStudy (anonymous):

Sorry I was attempting to answer it without having to write it on paper lol :)

OpenStudy (anonymous):

its okay! haha do you know how to do the second? what's the difference between varies directly and varies inversely?

OpenStudy (anonymous):

\[x = \frac{ k }{ y^2 }\] \[100\times6 = \frac{ k }{ 100 }\times 100 \] k = 600\[x = \frac{ 600 }{ 4 } or 150\]

OpenStudy (anonymous):

so basically inversely, it's divison and directly is multiplication?

OpenStudy (anonymous):

yeah pretty much

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

welcome